# SemEval-2024 Task 6: SHROOM, a Shared-task on Hallucinations and Related Observable Overgeneration Mistakes

### v2

@@ 4 Foraging and harvesting season: Collected data
All SHROOM data (models, outputs and annotations) are available under a CC-BY license. [^2] Link to be added.
@@ Appendix A Shared consciousnesses: Overview of approaches used by SHROOM teams
SHROOM datasets; Vectara model.
Fine-tuned models and voting classifier.
Alejandro Mosquera
SHROOM datasets; COMET, Vectara, LaBSE, GPT35 and GPT4 models.
Ensemble of publicly available models. Logistic Regression was used as final scoring model.
@@
SHROOM dataset, SBERT
Fully-connected neural network classifiers with SBERT embeddings as input.
AMEX AI LABS
SHROOM datasets; Vectara and OpenChat models.
Ensemble of LLM (using Openchat) zero shot and few shot with Vectara cross encoder based scores.
Atresa
(No report)
@@
LLaMA model.
Prompt-based approach with LLaMA.
BruceW
(No report)
@@
SHROOM dataset, data augmentation, RoBERTa
Finetuned a BERT or RoBERTa model with a softmax layer to output the probability of hallucinated text. Finetuning data is the labelled SHROOM data augmented with data points constructed by replacing words with synonyms.
CAISA
(No report)
@@
HalluEval dataset; Mistral 7B instruct model.
Ensemble of several role-based LLMs, which were either fine-tuned on hallucination data or role-based prompting.
daixiang
(No report)
deema
(No report)
@@
SHROOM dataset; OpenChat, DeBERTa, RoBERTa and T5 models.
Ensemble of several pretrained Transformer-based models to get features for validation and test data of SHROOM dataset and trained a boosting-based meta-model on top.
@@
SHROOM, LaBSE, T5, DistilUSE
Using pre-trained LLMs and classifiers
@@
SHROOM dataset, SBERT
Logistic regression and feed-forward classifier trained on SBERT embeddings
@@
SHROOM, labeled 3000 samples of the training data
Fine-tuned a DeBERTa-v3-large
@@
SHROOM datasets; GPT, SelfCheckGPT and Vectara models.
Prompts and GroupCheckGPT. NB: due to a team name change, this team is also referred to as GroupCheckGPT by some participants.
@@
SHROOM, SNLI, MNLI and PAWS datasets; Vectara and GPT4 models.
Highest results obtained with zero-shot prompting in the model-aware track; pretraining on NLI and PAWS followed by finetuning on the model-agnostic track.
@@
SHROOM with training dataset labeled using GPT-4; DeBERTaV3, InternLM2, SBERT, and UniEval.
Fine-tune the DeBERTaV3 and InternLM2 models, and call the SBERT and UniEval models to select the optimal threshold usinf SHROOM & syntheticaly labeled data. The system obtians the final results by combining the prediction results of each model.
@@
SHROOM datasets; Sentence-t5, BGE, e5 models.
Computes the cosine similarity of sentence embeddings and classify based on an empirical threshold value.
@@
DeBERTa models.
Zero shot inference, pretrained cross encoder model
@@
SHROOM model-agnostic dataset, DeBERTa pretrained and finetuned on MNLI, SOLAR-10.7B quantized from TheBloke (for synthetic data generation)
Encoder and classifier, fine-tuned in various ways (including with synthetic data)
@@
SHROOM dataset, SBERT, bart-large-mnli, Mixtral
Three approaches: (1) Cosine similarity of SBERT embeddings between source-hypothesis and source-target pairs; (2) NLI classification using bart-large-mnli model; and (3) Mixtral prompting. Only the Mixtral results were submitted.
@@
SHROOM dataset; Mixtral and RoBERTa models.
Mixtral prompting and RoBERTa finetuning.
@@
SHROOM, GPT-3.5, Sentence Transformers
Tried two approaches: (1) hypothesis-target cosine similarity, using a threshold value to determine whether the hypothesis is a hallucation. (2) SelfCheckGPT with a customized prompt for each NLG task, designed to assess its coherence with the provided source and target. Each prompt is iterated through the GPT-3.5 model five times, and the final label is determined by the majority response.
@@
RoBERTa
Used a pretrained model (roberta-large-openai-detector) that has been trained to distinguish between text generated by LLMs and text written by humans.
@@
SHROOM, Mistral-7B-Instruct-v0.2, self constructed training data
Supervised fine-tuning over synthetically constructed weakly supervised training data.
@@
Mistral2, LLaMa2, Phi2 and Zephyr models; uses SHROOM train set for prompt optimization.
Ensembling over the output logits of prompt-based LLMs (mistral, llama etc) after automatically optimizing their prompts ("OPRO").
@@
SHROOM dataset; GPT 3.5 and GPT 4 models.
In-context learning with role-play and automatic prompt generation in a few-shot classfier, using a closed-source LLM.
SibNN
SHROOM datasets; XLM-RoBERTa model.
Fine-tunes a self-adaptive hierarchical variant of XLM-RoBERTa-XL twice: first as an embedder (in a few-shot mode), then as a binary classifier. More details at
@@
silk_road
SHROOM datasets; Vectara model.
Fine-tunes an off-the-shelf Cross-Encoder hallucination evaluation model.
Skoltech
(No report)
@@
SHROOM datasets; LaBSE, DeBERTa, Zephyr, Mistral and Llama2 models.
Using two LLMs to classify and explain their decision and another LLM to judge and decide based on those explanations.
@@
SHROOM (synthetically augmented), QQP and PAWS datasets; E5, T5, Vectara models.
Fine-tuning of e5-mistral-7b-instruct using synthetic data collected with LLaMA2-7B adapters trained to produce data with and without hallucinations. However, there are two other systems: one works as a voting ensemble of multiple LLMs, and another uses the Mutual Implication Score architecture.
Team CentreBack
SHROOM dataset; DeBERTa model.
Uses an off-the-shelf library (SelfCheckGPT’s SelfCheckNLI function) to calculate contradiction scores on a small labeled test set and then defined a threshold for hallucination.
@@
SHROOM dataset; Vectara model.
Model-aware track best submissions uses a Vectara hallucination detection model finetuned on the validation set. The best model-agnostic track submission is a meta-model that utilizes linear regression and is trained on features that correspond to probabilities predicted by individual systems we implemented.
UCC-NLP
SHROOM dataset; GPT-3.5 and Vectara models.
Uses BertScore and GPT-3.5 to create synthetic labels and fine-tune a Vectara LLM.
@@
SHROOM dataset; TULU-DPO model.
Zero-shot approach
ustc_xsong
(No report)
zhuming
(No report)
@@
Mistral, Mixtral, LLaMA, Falcon, WizardLM and Capybara models.
Zero-shot prompt engineering. Expects most LLMs will have different hallucination patterns, and tests whether ensembling can mitigate this.

### v3

@@ 4 Foraging and harvesting season: Collected data
All SHROOM data (models, outputs and annotations) are available under a CC-BY license. [^2] See helsinki-nlp.github.io/shroom
@@ Appendix A Shared consciousnesses: Overview of approaches used by SHROOM teams
SHROOM datasets; Vectara model. Fine-tuned models and voting classifier.
Alejandro Mosquera SHROOM datasets; COMET, Vectara, LaBSE, GPT35 and GPT4 models. Ensemble of publicly available models. Logistic Regression was used as final scoring model.
@@
SHROOM dataset, SBERT Fully-connected neural network classifiers with SBERT embeddings as input.
AMEX AI LABS SHROOM datasets; Vectara and OpenChat models. Ensemble of LLM (using Openchat) zero shot and few shot with Vectara cross encoder based scores.
Atresa (No report)
@@
LLaMA model. Prompt-based approach with LLaMA.
BruceW (No report)
@@
SHROOM dataset, data augmentation, RoBERTa Finetuned a BERT or RoBERTa model with a softmax layer to output the probability of hallucinated text. Finetuning data is the labelled SHROOM data augmented with data points constructed by replacing words with synonyms.
CAISA (No report)
@@
HalluEval dataset; Mistral 7B instruct model. Ensemble of several role-based LLMs, which were either fine-tuned on hallucination data or role-based prompting.
daixiang (No report)
deema (No report)
@@
SHROOM dataset; OpenChat, DeBERTa, RoBERTa and T5 models. Ensemble of several pretrained Transformer-based models to get features for validation and test data of SHROOM dataset and trained a boosting-based meta-model on top.
@@
SHROOM, LaBSE, T5, DistilUSE Using pre-trained LLMs and classifiers
@@
SHROOM dataset, SBERT Logistic regression and feed-forward classifier trained on SBERT embeddings
@@
SHROOM, labeled 3000 samples of the training data Fine-tuned a DeBERTa-v3-large
@@
SHROOM datasets; GPT, SelfCheckGPT and Vectara models. Prompts and GroupCheckGPT. NB: due to a team name change, this team is also referred to as GroupCheckGPT by some participants.
@@
SHROOM, SNLI, MNLI and PAWS datasets; Vectara and GPT4 models. Highest results obtained with zero-shot prompting in the model-aware track; pretraining on NLI and PAWS followed by finetuning on the model-agnostic track.
@@
SHROOM with training dataset labeled using GPT-4; DeBERTaV3, InternLM2, SBERT, and UniEval. Fine-tune the DeBERTaV3 and InternLM2 models, and call the SBERT and UniEval models to select the optimal threshold usinf SHROOM & syntheticaly labeled data. The system obtians the final results by combining the prediction results of each model.
@@
SHROOM datasets; Sentence-t5, BGE, e5 models. Computes the cosine similarity of sentence embeddings and classify based on an empirical threshold value.
@@
DeBERTa models. Zero shot inference, pretrained cross encoder model
@@
SHROOM model-agnostic dataset, DeBERTa pretrained and finetuned on MNLI, SOLAR-10.7B quantized from TheBloke (for synthetic data generation) Encoder and classifier, fine-tuned in various ways (including with synthetic data)
@@
SHROOM dataset, SBERT, bart-large-mnli, Mixtral Three approaches: (1) Cosine similarity of SBERT embeddings between source-hypothesis and source-target pairs; (2) NLI classification using bart-large-mnli model; and (3) Mixtral prompting. Only the Mixtral results were submitted.
@@
SHROOM dataset; Mixtral and RoBERTa models. Mixtral prompting and RoBERTa finetuning.
@@
SHROOM, GPT-3.5, Sentence Transformers Tried two approaches: (1) hypothesis-target cosine similarity, using a threshold value to determine whether the hypothesis is a hallucation. (2) SelfCheckGPT with a customized prompt for each NLG task, designed to assess its coherence with the provided source and target. Each prompt is iterated through the GPT-3.5 model five times, and the final label is determined by the majority response.
@@
RoBERTa Used a pretrained model (roberta-large-openai-detector) that has been trained to distinguish between text generated by LLMs and text written by humans.
@@
SHROOM, Mistral-7B-Instruct-v0.2, self constructed training data Supervised fine-tuning over synthetically constructed weakly supervised training data.
@@
Mistral2, LLaMa2, Phi2 and Zephyr models; uses SHROOM train set for prompt optimization. Ensembling over the output logits of prompt-based LLMs (mistral, llama etc) after automatically optimizing their prompts ("OPRO").
@@
SHROOM dataset; GPT 3.5 and GPT 4 models. In-context learning with role-play and automatic prompt generation in a few-shot classfier, using a closed-source LLM.
SibNN SHROOM datasets; XLM-RoBERTa model. Fine-tunes a self-adaptive hierarchical variant of XLM-RoBERTa-XL twice: first as an embedder (in a few-shot mode), then as a binary classifier. More details at
@@
silk_road SHROOM datasets; Vectara model. Fine-tunes an off-the-shelf Cross-Encoder hallucination evaluation model.
Skoltech (No report)
@@
SHROOM datasets; LaBSE, DeBERTa, Zephyr, Mistral and Llama2 models. Using two LLMs to classify and explain their decision and another LLM to judge and decide based on those explanations.
@@
SHROOM (synthetically augmented), QQP and PAWS datasets; E5, T5, Vectara models. Fine-tuning of e5-mistral-7b-instruct using synthetic data collected with LLaMA2-7B adapters trained to produce data with and without hallucinations. However, there are two other systems: one works as a voting ensemble of multiple LLMs, and another uses the Mutual Implication Score architecture.
Team CentreBack SHROOM dataset; DeBERTa model. Uses an off-the-shelf library (SelfCheckGPT’s SelfCheckNLI function) to calculate contradiction scores on a small labeled test set and then defined a threshold for hallucination.
@@
SHROOM dataset; Vectara model. Model-aware track best submissions uses a Vectara hallucination detection model finetuned on the validation set. The best model-agnostic track submission is a meta-model that utilizes linear regression and is trained on features that correspond to probabilities predicted by individual systems we implemented.
UCC-NLP SHROOM dataset; GPT-3.5 and Vectara models. Uses BertScore and GPT-3.5 to create synthetic labels and fine-tune a Vectara LLM.
@@
SHROOM dataset; TULU-DPO model. Zero-shot approach
ustc_xsong (No report)
zhuming (No report)
@@
Mistral, Mixtral, LLaMA, Falcon, WizardLM and Capybara models. Zero-shot prompt engineering. Expects most LLMs will have different hallucination patterns, and tests whether ensembling can mitigate this.